OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x-9y=0 -x+3y=-3 A. 0 B. 1 C. Infinitely many

OpenStudy (anonymous):

3x-9y=0 -x+3y=-3 (3) ____________ 3x-9y=0 -3x+9y=-9 ______________ 0x+0y=-9 it doesn't have solutions.

OpenStudy (anonymous):

Thank you.